Problem 12. Find the equation of the line that passes through two points (4.57, 8.29) and (7.17, -2.69). Round the equation in slope-intercept form. Round the slope and the constant term to two decimal places
Answer #1

Here we are given the two points ( -4.57,8.29) and (7.17,-2.69)

So x y

-4.57 8.29

7.17 -2.69

So the slope =(y2-y1)/(x2-x1) = -10.98/11.74

Slope = -0.94

The intercept of line is

Intercept = ybar - b1 * x bar

Intercept = 2.8 - (-0.94)*1.3

Intercept = 4.02

The equation of line, y = 4.02 - 0.94 * x